Decision on Accepting Additional Business

Homestead Jeans Co. has an annual plant capacity of 63,800 units, and current production is 45,600 units. Monthly fixed costs are $41,300, and variable costs are $25 per unit. The present selling price is $37 per unit. On November 12 of the current year, the company received an offer from Dawkins Company for 16,600 units of the product at $28 each. Dawkins Company will market the units in a foreign country under its own brand name. The additional business is not expected to affect the domestic selling price or quantity of sales of Homestead Jeans Co.

a. Prepare a differential analysis dated November 12 on whether to reject (Alternative 1) or accept (Alternative 2) the Dawkins order. If an amount is zero, enter "0". For those boxes in which you must enter subtracted or negative numbers use a minus sign.

Differential Analysis
Reject Order (Alt. 1) or Accept Order (Alt. 2)
November 12
Reject Order (Alternative 1) Accept Order (Alternative 2) Differential Effect on Income (Alternative 2)
Revenues $ $ $
Costs:
Variable manufacturing costs
Income (Loss) $ $ $

b. Having unused capacity available is   to this decision. The differential revenue is   than the differential cost. Thus, accepting this additional business will result in a net  .

c. What is the minimum price per unit that would produce a positive contribution margin? Round your answer to two decimal places.
$
Answer #1

a. Prepare a differential analysis dated November 12 on whether to reject (Alternative 1) or accept (Alternative 2) the Dawkins order. If an amount is zero, enter "0". For those boxes in which you must enter subtracted or negative numbers use a minus sign.

Differential Analysis
Reject Order (Alt. 1) or Accept Order (Alt. 2)
November 12
Reject Order (Alternative 1) Accept Order (Alternative 2) Differential Effect on Income (Alternative 2)
Revenues $0 16600*28 = 464800 $464800
Costs:
Variable manufacturing costs 0 16600*25 = -415000 -415000
Income (Loss) $0 49800 49800

b) Having unused capacity available is accept the special order to this decision. The differential revenue is   more than the differential cost. Thus, accepting this additional business will result in a net income 49800

c) Minimum price per unit = 26 per unit